# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 7.1.2 Ferromagnetic systems
Many works have considered the Markovian framework for anomalous relaxations in anti-ferromagnetic systems, and a rich phase diagram of anomalous effects was demonstrated in many variants of anti-ferromagnetic systems, e.g., [3, 65, 96, 82, 184]. It is therefore natural to consider also the case of a ferromagnetic systems. Recently, a series of investigations [73, 185, 186, 187] reports numerical observations of a different type of anomalous effect in ferromagnetic systems. The ferromagnetic systems surveyed include the 2D Ising model [73], the q-Potts model in 2D and with varying interaction lengths [185], and the 3D Ising model along with the 2D Ising model with long-range order [187]. In these numerical results, the system is initiated at equilibrium corresponding to a temperature above the phase transition (but with the specific constraint that half the spins are up and half down  [73, 186]). After the preparation, the system is instantaneously quenched to a temperature below the phase transition. The quenching means that the system starts to follow a single spin flip dynamics, with Glauber or Metropolis flipping rates, where the temperature corresponds to the bath temperature. Rather than looking for the phase transition point, the energy per spin is monitored as a function of the simulation time. It is numerically observed that although the initially higher temperature system has a higher energy per spin, after long enough evolution with dynamic corresponding to bath temperature below the phase transition, the initially hotter system has a lower energy per spin. In addition to observing the energy per spin, in [73] crossing between average domain length was also observed.
Interestingly, the exact nature of the observed effect is not understood at this point. For example, it seems to be important for the observed effect that the initial condition have an exactly equal mix of up and down spins (and not the Boltzmann distribution of up and down spins expected in equilibrium), but the reason is unclear. Note that even though above the phase transition point the mean magnetization is expected to be zero in average, the variance scales as the square root of the number of spins, hence having exactly the same number of up and down spins is a very strong constraint. It is quite possible that this effect is related to the Markovian Mpemba effect discussed in Sec. 2. Unfortunately, finding the slowest decaying mode in a large 2D Ising model (ref. [73] used an array of $256\times 256$ spins) is not possible analytically or even numerically, and no trace for Markovian Mpemba effect was found for smaller systems where analytic calculations are possible. One route that might determine the existence of a Mpemba effect without the need to evaluate $a_{2}$ could, in principle, come from the study of the relaxation of relevant observables, as outlined in the following section as well as in Sec. 2.1.6.
@@ 7.2.1 Molecular dynamics simulations
Four works [47, 48, 143, 188] used molecular dynamics simulations to try to shed light on the Mpemba effect in water. The first, [47], considered molecular dynamics simulations of up to 1000 molecules. The initial configurations were sampled from equilibrium distributions corresponding to the temperature before the quench (from 280 K to 390 K). Three different force field models were used (Molinero with a Stillinger-Weber angular function, SPC rigid water model and the F3C flexible water model), all in thermostated NPT simulation (Nose-Hoover thermostat), to a final temperature of 100 K. In each case, the vibrational density of states (DOS) was calculated, and the population of states in the region 80–160 cm-1, which is known to be important for water crystallization, was recorded. It was found that quenching from warm water (above 370 K) led to a higher population in the relevant density of states at this specific energy interval, in comparison to ice. By contrast, a system quenched from cool water (below 300 K) has a lower population in this energy range. The observations might suggest that the relaxation of vibrational states to that of ice is faster in the initially hot system. This work also explored the importance of the specific range of DOS and found that the specific regime is associated with internal vibrations of water hexagons or perhaps fused decagons.
@@
Figure 37: Numerical results of the TIP4P/Ice model with 96 molecules. The upper-left figure shows the potential energy as a function of time for various initial temperatures, all quenched to $T_{f}=230\,{\rm K}$. The symbols show the locations of the jump based on which the phase transition time is declared (with additional symbols for other trajectories that are not plotted). The initial part of the relaxation is shown in the upper right figure. The lower figure shows that the mean freezing time is a nonmonotonic function of the initial temperature for the chosen set of parameters. The figure is reprinted from  [188].
@@ 7.3 Simulations in kinetic theory
The direct simulation Monte Carlo (DSMC) method, due originally to Bird [192], solves the Boltzmann equation and can be applied to simulate rarefied gas dynamics. The basic ideas can be adapted to a wide range of models of dilute granular gases (see Sec. 4). The space region is divided into a series of cells whose dimension must be smaller than the mean free path, so that flows are approximately constant inside the cell cavities. The flow of particles between collisions is considered uncoupled over a time interval $\Delta t$, which must be smaller than the average time between collisions. Next, each particle is moved a distance $v_{i}\Delta t$, where $v_{i}$ is the velocity of the $i$-th particle. Then the collision part of the algorithm is resolved. When the collision process is computed, only particles inside the same cell can collide; otherwise, the positions of particles play no role. A crucial aspect is that, in each cell, pairs of particles are chosen randomly, and their collision is accepted or rejected according to a probability that is proportional to the relative velocity between them. If the collision is accepted, the post-collisional velocities are computed using an impact parameter chosen at random from a uniform distribution of incident particles. The time counter is advanced after collisions and the process repeated in each cavity until its time counter exceeds the time step $\Delta t$ [193]. The DSMC method is a faster, more efficient way to simulate the dilute granular dynamics. See Figs. 22 and 23 for a comparison between methods and theoretical predictions.
@@ 8.1.1 Optimal heating and cooling protocols
A natural application for many variants of the Mpemba effect is optimal cooling (or heating in the case of an inverse Mpemba effect). If cooling a system initiated at $T_{\mathrm{hot}}$ is faster than cooling the same system initiated at $T_{\mathrm{medium}}$ when the cooling mechanism is coupling the system to a bath at $T_{\mathrm{cold}}$, then perhaps cooling a system initiated at $T_{\mathrm{medium}}$ by first heating it towards $T_{\mathrm{hot}}$ and then coupling it to $T_{\mathrm{cold}}$ is faster than coupling it to a thermal bath at $T_{\mathrm{cold}}$ without the pre-heating stage. But if pre-heating takes longer than the time gained during the cooling stage, the extra step will not reduce the overall protocol time.
@@
In most cases, finding the exact optimal protocol is very difficult, even numerically. However, for systems with only a few states, this is sometimes possible: in optimal-control theory, finding $T_{b}(t)$ that minimizes $a_{n+1}$ subject to $a_{2}=a_{3}=\ldots=a_{n}=0$ is a quadratic end-point minimization problem with linear end-point constraints, a bounded domain and no path cost. Such minimization problems can be addressed numerically using Pontryagin’s maximum principle [196]. Since there is no path cost, they can also be solved numerically with methods such as sequential quadratic programming [197]. In [65], an example for optimal heating in a four-state system where at a given finite time $a_{2}=a_{3}=0$ and $|a_{4}|$ is minimal was demonstrated. Unfortunately, when the number of states become large, this optimization problem becomes intractable. Nevertheless, it is often possible to find a protocol that sets $a_{2}=0$ or even $a_{2}=a_{3}=0$ at a finite time. These protocols are not expected to be optimal, but they are superior to the naive protocol of coupling only to the final temperature. Examples for such protocols were given for a Brownian particle diffusing in a potential, and for the 2D nearest-neighbor-coupling anti-ferromagnetic Ising model.
@@
Figure 38: Improving heating by first cooling the system (figure taken from [65]). (a) The system is a Brownian particle diffusing in the plotted 1D potential and reflecting boundary conditions. (b) A plot of $a_{2}$ as a function of temperature for $T<T_{\mathrm{hot}}$. The final temperature $T_{\mathrm{hot}}$ is represented by the red square, the initial temperature $T_{\mathrm{init}}$ by the black circle, and the cold temperature $T_{\mathrm{cold}}$ to which the system is coupled in the pre-cooling state by the blue triangle. The plot of $a_{2}$ demonstrates that there is no inverse Mpemba effect, weak or strong, for $T_{\mathrm{hot}}$. (c) The Kullback-Leibler divergence of the distribution $p(t)$ from the equilibrium at $T_{\mathrm{hot}}$ for the naive heating protocol where the system is coupled to $T_{\mathrm{hot}}$ (dashed orange line), and the pre-cooling protocol, where the system is first coupled to $T_{\mathrm{cold}}$ for some finite time, and only then it is coupled to $T_{\mathrm{hot}}$ (solid green line). The figures are reproduced from [65].
@@ 8.1.2 Optimal transport and heat engine efficiency
Subsequently, the authors examined discrete cases of Markov jump processes, where the rates obey detailed balance and are parameterized by energies $\{E_{j}\}$ and barriers $\{B_{i\!j}\}$ such that $W_{i\!j}\propto\exp\left(-(E_{j}-B_{i\!j})/T_{b}\right)$, as noted in  Eq. (17). In these discrete cases, they identify and analyze scenarios where the same dynamics produces a strong Mpemba effect and minimal total entropy dissipation. It would be intriguing to generalize these findings on optimal transport and the Mpemba effect to larger networks or more complex physical systems and to determine the aspects of relaxation protocols that simultaneously yield a Mpemba effect and minimal total entropy dissipation.
@@
$W_{21}(T_{b})\propto e^{-(E_{2}-E_{1})/T_{b}},\quad W_{13}(T_{b})\propto e^{-(E_{1}-E_{3})/T_{b}},\quad\text{and}\quad W_{32}(T_{b},\delta)\propto e^{-(E_{3}-E_{2})\delta/T_{b}}.$ (138)
@@
Figure 40: The device (mechanical Maxwell’s demon here) interacts with tape and work reservoir. The device and the tape are modeled as a three-level Markov jump system. They form a coupled system during the working cycle. b) Kullback-Leibler divergence as a function time for the same initial conditions and three different load distribution factors. The system has a Strong Mpemba effect at load factor $\delta_{\rm SM}=0.35$ for parameters: $T_{t}=0.824$, $T_{b}=2$, $\epsilon_{t}=0.9$ and $\epsilon_{d}=0.4$. c) Cycle time, $\tau_{\rm cyc}$ is chosen as the time where the KL divergence has a cutoff (here $10^{-12}$). The cycle time has a minimum at $\delta_{\rm SM}=0.35$, where we also have the Strong Mpemba effect. The figures are adapted from [98].
Closely related are work heat engine, that played an important role in the development of thermodynamic. A simple heat engine model where a three-state Markov jump process in which the temperature is periodically switched between two temperature , and thus it serves as the “working fluid” of a cyclic Otto engine, was studied in [207]. In this work, the authors compares the evolution of the system from the same initial conditions, both with and without the Mpemba effect. They demonstrate that the Mpemba effect can enhance the machine’s performance by decreasing the time required to reach the steady state. This shorter cycle results in an increased power output for a given efficiency, while maintaining stability. These findings align with previous results in  [98] and [200].
@@ 8.1.4 Rate of chemical reactions
where $k$ is the rate constant for the elemental reaction, $[E]$ and $[S]$ the enzyme and substrate concentrations, and $K$ the dissociation constant between enzyme and substrate. The Michaelis-Menten kinetics of Eq. (139) suggests that increasing enzyme concentrations typically accelerate enzymatic reactions. Hatakeyama considers a biomolecule with multisite modifications catalyzed by enzymes. For a single enzyme, the conventional Michaelis-Menten scenario holds; however, for two enzymes, an increase in the concentration of one can slow relaxation to a steady state, even if both enzymes individually act as catalysts. As a paradigm, Hatakeyama uses an allosteric dimer, where each monomer represents a modification site, and where monomers can form two different structures. The equilibrium distribution is independent of the number of enzymes, but the relaxation times depend on the population of each particular enzyme. The two enzymes catalyze reactions by changing the modification site. They show a branched network structure, with one of the branches having a local energy minimum (a metastable state). The substrates in the metastable state sequestrate the enzymes; these molecules confiscate all enzymes, and the system stays in the metastable state for a long time. In this case, two enzymes have different affinities for substrates, and one preferentially binds to a substrate. Thus, together they increase the flux to the metastable state and slowing down the relaxation toward the target state. The role of temperature here is replaced by the enzyme concentration. Hatakeyama’s work shows an example of a system where the relaxation of the system toward a steady state is a nonmonotonic function of the enzyme concentration. It was also noted int this work that two or more enzymes of different affinities for substrates are needed for the result.
@@
Figure 41: Phase space of a fully connected three-level system characterized by energies $E_{1}=0$, $E_{2}$, and $E_{3}$ and situated within a thermal bath at a temperature $T_{b}=1$. The load distribution factor is set to $\delta=0$, as introduced in  Eq. (140). In this space, the gray regions indicate areas where there is no strong Mpemba effect. The red regions represent areas with a strong Mpemba effect during heating, which have a parity index given by $\mathcal{P}\left(\mathcal{I}_{M}^{\textrm{inv}}\right)=1$. Conversely, the blue regions indicate the existence of the strong Mpemba effect during cooling, with a parity index described by $\mathcal{P}\left(\mathcal{I}_{M}^{\textrm{dir}}\right)=1$. The definitions of the parity indices can be found in Eqs. (41) and (42). Moreover, the red and blue regions have the highest rate (the global maximum of rates) between the highest and lowest energy levels. The swirling cyan arrow illustrates the chirality of the arms associated with the strong Mpemba effect. The figure is adapted from [98].
@@
The Mpemba effect in linear reaction networks, which has the same flavor as the aobe mntioned work, was studied in  [98]. The authors show examples where changing the dynamics allows one to choose, for a given initial condition, dynamics with anomalous thermal relaxation. They provide analytical results and insights regarding when the strong Mpemba effect occurs in unimolecular reactions involving three and four species, depending on the dynamics. To quantify these dynamics, the authors introduce a load-distribution factor, $\delta$, which modifies the forward and backward rates while maintaining detailed balance. Specifically, the transition rates are defined as
$W_{i\!j}=\Gamma e^{-(E_{j}-E_{i})\delta/T_{b}}\quad\text{and}\quad W_{\!ji}=\Gamma e^{-(E_{j}-E_{i})(1-\delta)/T_{b}},$ (140)

### v3

@@ 7.1.2 Ferromagnetic systems
Many works have considered the Markovian framework for anomalous relaxations in antiferromagnetic systems, and a rich phase diagram of anomalous effects was demonstrated in many variants of antiferromagnetic systems, e.g., [3, 65, 96, 82, 193]. It is therefore natural to consider also the case of a ferromagnetic systems. Recently, a series of investigations [73, 194, 195, 196] reports numerical observations of a different type of anomalous effect in ferromagnetic systems. The ferromagnetic systems surveyed include the 2D Ising model [73], the q-Potts model in 2D and with varying interaction lengths [194], and the 3D Ising model along with the 2D Ising model with long-range order [196]. In these numerical results, the system is initiated at equilibrium corresponding to a temperature above the phase transition (but with the specific constraint that half the spins are up and half down  [73, 195]). After the preparation, the system is instantaneously quenched to a temperature below the phase transition. The quenching means that the system starts to follow a single spin flip dynamics, with Glauber or Metropolis flipping rates, where the temperature corresponds to the bath temperature. Rather than looking for the phase transition point, the energy per spin is monitored as a function of the simulation time. It is numerically observed that although the initially higher temperature system has a higher energy per spin, after long enough evolution with dynamic corresponding to bath temperature below the phase transition, the initially hotter system has a lower energy per spin. In addition to observing the energy per spin, in [73] crossing between average domain length was also observed.
Interestingly, the exact nature of the observed effect is not understood at this point. For example, it seems to be important for the observed effect that the initial condition have an exactly equal mix of up and down spins (and not the Boltzmann distribution of up and down spins expected in equilibrium), but the reason is unclear. Note that even though above the phase transition point the mean magnetization is expected to be zero on average, the variance scales as the square root of the number of spins; hence, having exactly the same number of up and down spins is a very strong constraint. It is quite possible that this effect is related to the Markovian Mpemba effect discussed in Sec. 2. Unfortunately, finding the slowest decaying mode in a large 2D Ising model (ref. [73] used an array of $256\times 256$ spins) is not possible analytically or even numerically, and no trace for Markovian Mpemba effect was found for smaller systems where analytic calculations are possible. One route that might determine the existence of a Mpemba effect without the need to evaluate $a_{2}$ could, in principle, come from the study of the relaxation of relevant observables, as outlined in the following section as well as in Sec. 2.1.6.
@@ 7.2.1 Molecular dynamics simulations
Four works [47, 48, 150, 197] used molecular dynamics simulations to try to shed light on the Mpemba effect in water. The first, [47], considered molecular dynamics simulations of up to 1000 molecules. The initial configurations were sampled from equilibrium distributions corresponding to the temperature before the quench, from 280 K to 390 K. Three different force field models were used (Molinero with a Stillinger-Weber angular function, SPC rigid water model and the F3C flexible water model), all in thermostated NPT simulation (Nose-Hoover thermostat), to a final temperature of 100 K. In each case, the vibrational density of states (DOS) was calculated, and the population of states in the region 80–160 cm-1, which is known to be important for water crystallization, was recorded. It was found that quenching from warm water (above 370 K) led to a higher population in the relevant density of states at this specific energy interval, in comparison to ice. By contrast, a system quenched from cool water (below 300 K) has a lower population in this energy range. The observations might suggest that the relaxation of vibrational states to that of ice is faster in the initially hot system. This work also explored the importance of the specific range of DOS and found that the specific regime is associated with internal vibrations of water hexagons or perhaps fused decagons.
@@
Figure 37: Numerical results of the TIP4P/Ice model with 96 molecules. The upper-left figure shows the potential energy as a function of time for various initial temperatures, all quenched to $T_{f}=230\,{\rm K}$. The symbols show the locations of the jump based on which the phase transition time is declared (with additional symbols for other trajectories that are not plotted). The initial part of the relaxation is shown in the upper right figure. The lower figure shows that the mean freezing time is a nonmonotonic function of the initial temperature for the chosen set of parameters. Source: Reprinted with permission from [197].
@@ 7.3 Simulations in kinetic theory
The direct simulation Monte Carlo (DSMC) method, due originally to Bird [201], solves the Boltzmann equation and can be applied to simulate rarefied gas dynamics. The basic ideas can be adapted to a wide range of models of dilute granular gases; see Sec. 4. The space region is divided into a series of cells whose dimension must be smaller than the mean free path, so that flows are approximately constant inside the cell cavities. The flow of particles between collisions is considered uncoupled over a time interval $\Delta t$, which must be smaller than the average time between collisions. Next, each particle is moved a distance $v_{i}\Delta t$, where $v_{i}$ is the velocity of the $i$-th particle. Then the collision part of the algorithm is resolved. When the collision process is computed, only particles inside the same cell can collide; otherwise, the positions of particles play no role. A crucial aspect is that, in each cell, pairs of particles are chosen randomly, and their collision is accepted or rejected according to a probability that is proportional to the relative velocity between them. If the collision is accepted, the post-collisional velocities are computed using an impact parameter chosen at random from a uniform distribution of incident particles. The time counter is advanced after collisions and the process repeated in each cavity until its time counter exceeds the time step $\Delta t$ [202]. The DSMC method is a faster, more efficient way to simulate the dilute granular dynamics. See Figs. 22 and 23 for a comparison between methods and theoretical predictions.
@@ 8.1.1 Optimal heating and cooling protocols
A natural application for many variants of the Mpemba effect is optimal cooling, or heating in the case of an inverse Mpemba effect. If cooling a system initiated at $T_{\mathrm{hot}}$ is faster than cooling the same system initiated at $T_{\mathrm{medium}}$ when the cooling mechanism is coupling the system to a bath at $T_{\mathrm{cold}}$, then perhaps cooling a system initiated at $T_{\mathrm{medium}}$ by first heating it towards $T_{\mathrm{hot}}$ and then coupling it to $T_{\mathrm{cold}}$ is faster than coupling it to a thermal bath at $T_{\mathrm{cold}}$ without the pre-heating stage. But if pre-heating takes longer than the time gained during the cooling stage, the extra step will not reduce the overall protocol time.
@@
In most cases, finding the exact optimal protocol is very difficult, even numerically. However, for systems with only a few states, this is sometimes possible: in optimal-control theory, finding $T_{b}(t)$ that minimizes $a_{n+1}$ subject to $a_{2}=a_{3}=\ldots=a_{n}=0$ is a quadratic end-point minimization problem with linear end-point constraints, a bounded domain and no path cost. Such minimization problems can be addressed numerically using Pontryagin’s maximum principle [205]. Since there is no path cost, they can also be solved numerically with methods such as sequential quadratic programming [206]. In [65], an example for optimal heating in a four-state system where at a given finite time $a_{2}=a_{3}=0$ and $|a_{4}|$ is minimal was demonstrated. Unfortunately, when the number of states become large, this optimization problem becomes intractable. Nevertheless, it is often possible to find a protocol that sets $a_{2}=0$ or even $a_{2}=a_{3}=0$ at a finite time. These protocols are not expected to be optimal, but they are superior to the naive protocol of coupling only to the final temperature. Examples for such protocols were given for a Brownian particle diffusing in a potential, and for the 2D nearest-neighbor-coupling antiferromagnetic Ising model.
@@
Figure 38: Improving heating by first cooling the system (figure taken from [65]). (a) The system is a Brownian particle diffusing in the plotted 1D potential and reflecting boundary conditions. (b) A plot of $a_{2}$ as a function of temperature for $T<T_{\mathrm{hot}}$. The final temperature $T_{\mathrm{hot}}$ is represented by the red square, the initial temperature $T_{\mathrm{init}}$ by the black circle, and the cold temperature $T_{\mathrm{cold}}$ to which the system is coupled in the pre-cooling state by the blue triangle. The plot of $a_{2}$ demonstrates that there is no inverse Mpemba effect, weak or strong, for $T_{\mathrm{hot}}$. (c) The Kullback-Leibler divergence of the distribution $p(t)$ from the equilibrium at $T_{\mathrm{hot}}$ for the naive heating protocol where the system is coupled to $T_{\mathrm{hot}}$ (dashed orange line), and the pre-cooling protocol, where the system is first coupled to $T_{\mathrm{cold}}$ for some finite time, and only then it is coupled to $T_{\mathrm{hot}}$ (solid green line). Source: Reprinted with permission from [65].
@@ 8.1.2 Optimal transport and heat engine efficiency
Subsequently, the authors examined discrete cases of Markov jump processes, where the rates obey detailed balance and are parameterized by energies $\{E_{j}\}$ and barriers $\{B_{i\!j}\}$ such that $W_{i\!j}\propto\exp\left(-(E_{j}-B_{i\!j})/k_{B}T_{b}\right)$, as noted in  Eq. (17). In these discrete cases, they identify and analyze scenarios where the same dynamics produces a strong Mpemba effect and minimal total entropy dissipation. It would be intriguing to generalize these findings on optimal transport and the Mpemba effect to larger networks or more complex physical systems and to determine the aspects of relaxation protocols that simultaneously yield a Mpemba effect and minimal total entropy dissipation.
@@
$W_{21}(T_{b})\propto e^{-(E_{2}-E_{1})/k_{B}T_{b}},\quad W_{13}(T_{b})\propto e^{-(E_{1}-E_{3})/k_{B}T_{b}},\quad\text{and}\quad W_{32}(T_{b},\delta)\propto e^{-(E_{3}-E_{2})\delta/k_{B}T_{b}}.$ (138)
@@
Figure 40: Mpemba effect in a Maxwell demon. (a) The device (mechanical Maxwell’s demon here) interacts with tape and work reservoir. The device and the tape are modeled as a three-level Markov jump system. They form a coupled system during the working cycle. (b) Kullback-Leibler divergence as a function of time for the same initial conditions and three different load distribution factors. The system has a strong Mpemba effect at load factor $\delta_{\rm SM}=0.35$ for parameters: $T_{t}=0.824$, $T_{b}=2$, $\epsilon_{t}=0.9$ and $\epsilon_{d}=0.4$. (c) Cycle time, $\tau_{\rm cyc}$ is chosen as the time when the KL divergence has a cutoff (here $10^{-12}$). The cycle time has a minimum at $\delta_{\rm SM}=0.35$, where we also have the strong Mpemba effect. Source: Adapted with permission from [97].
Closely related are work-heat engines, which played an important role in the development of thermodynamics. A simple heat-engine model where a three-state Markov jump process in which the temperature is periodically switched between two temperatures. It thus serves as the “working fluid” of a cyclic Otto engine [216]. In this work, the authors compare the evolution of the system from the same initial conditions, both with and without the Mpemba effect. They demonstrate that the Mpemba effect can enhance the machine’s performance by decreasing the time required to reach steady state. The shorter cycle increases power output for a given efficiency, while maintaining stability. The findings align with previous results in  [97] and [209]. An analogous quantum version of the Otto engine with a quantum Mpemba effect was studied in [217].
@@ 8.1.4 Rate of chemical reactions
where $k$ is the rate constant for the elemental reaction, $[E]$ and $[S]$ the enzyme and substrate concentrations, and $K$ the dissociation constant between enzyme and substrate. The Michaelis-Menten kinetics of Eq. (139) suggests that increasing enzyme concentrations typically accelerate enzymatic reactions. Hatakeyama considers a biomolecule with multisite modifications catalyzed by enzymes. For a single enzyme, the conventional Michaelis-Menten scenario holds; however, for two enzymes, an increase in the concentration of one can slow relaxation to a steady state, even if both enzymes individually act as catalysts. As a paradigm, Hatakeyama uses an allosteric dimer, where each monomer represents a modification site, and where monomers can form two different structures. The equilibrium distribution is independent of the number of enzymes, but the relaxation times depend on the population of each particular enzyme. The two enzymes catalyze reactions by changing the modification site. They show a branched network structure, with one of the branches having a local energy minimum (a metastable state). The substrates in the metastable state sequestrate the enzymes; these molecules remove all enzymes, and the system stays in the metastable state for a long time. In this case, two enzymes have different affinities for substrates, and one preferentially binds to a substrate. Together, they increase the flux to the metastable state and slow the relaxation toward the target state. The role of temperature here is replaced by the enzyme concentration. Hatakeyama’s work shows an example of a system where the relaxation of the system toward a steady state is a nonmonotonic function of the enzyme concentration. It was also noted in this work that two or more enzymes of different affinities for substrates are needed for the result.
@@
Figure 41: Phase space of a fully connected three-level system characterized by energies $E_{1}=0$, $E_{2}$, and $E_{3}$ and situated within a thermal bath at a temperature $T_{b}=1$. The load distribution factor is set to $\delta=0$, as introduced in  Eq. (140). In this space, the gray regions indicate areas where there is no strong Mpemba effect. The red regions represent areas with a strong Mpemba effect during heating, which have a parity index given by $\mathcal{P}\left(\mathcal{I}_{M}^{\textrm{inv}}\right)=1$. Conversely, the blue regions indicate the existence of the strong Mpemba effect during cooling, with a parity index described by $\mathcal{P}\left(\mathcal{I}_{M}^{\textrm{dir}}\right)=1$. The definitions of the parity indices can be found in Eqs. (41) and (42). Moreover, the red and blue regions have the highest rate (the global maximum of rates) between the highest and lowest energy levels. The swirling cyan arrow illustrates the chirality of the arms associated with the strong Mpemba effect. Source: Adapted with permission from [97].
@@
A Mpemba effect in linear reaction networks, which has the same flavor as the above work, was studied in  [97]. The authors show examples where changing the dynamics allows one to choose, for a given initial condition, dynamics with anomalous thermal relaxation. They provide analytical results and insights regarding when the strong Mpemba effect occurs in unimolecular reactions involving three and four species, depending on the dynamics. To quantify these dynamics, the authors introduce a load-distribution factor, $\delta$, which modifies the forward and backward rates while maintaining detailed balance. Specifically, the transition rates are defined as
$W_{i\!j}=\Gamma e^{-(E_{j}-E_{i})\delta/k_{B}T_{b}}\quad\text{and}\quad W_{\!ji}=\Gamma e^{-(E_{j}-E_{i})(1-\delta)/k_{B}T_{b}},$ (140)